# Causal Inference in Longitudinal Data under Unknown Interference

### v4

@@ A.4 Proof for Theorem 2
Kojevnikov, Marmer and Song (2021) show that $\tilde{\Sigma}_{Y}(d)\to\Sigma_{Y}(d)$ as $N\to\infty$. Let $\mathbf{R}(d)=\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yi}(d)\right]{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yj}^{\prime}(d)\right]$. Then, by Lemma A.2,
@@
A.5 Technical Details of the Extensions
Placebo Tests
Figure A.1: A DAG illustration for placebo tests
[…]
Dependence in Treatment Assignment
Assumption A.1 (Design mixing and design separation)
$\alpha_{INT,t}=o(N),\alpha_{EXT,t}=o(N),$
for any period $t$.
This assumption is satisfied by many common treatment assignment mechanisms, such as complete randomization. Formal proofs can be found in Sävje, Aronow and Hudgens (2021).
Marginal Structural Models with a Large $T$
Appendix B Extra Evidence from Simulation and Applications
B.1 Examples of $d_{ij}$ and $\Omega_{j}(d)$
Table B.1: Examples of $d_{ij}$ and $\Omega_{j}(d)$
@@ B.1 Examples of $d_{ij}$ and $\Omega_{j}(d)$
[…]
@@
B.2 Simulation Evidence
In this section, we test the performance of the proposed method using the simulated dataset examined in the main text, where $N=400$ and $T=2$. We focus on the setting in which the units are embedded in a social network (top-left plot of Figure 2), generated by a random geometric graph (RGG) model (Penrose, 2003; Leung, 2022). Let $G_{ij}$ indicate the presence of an edge between units $i$ and $j$. Then $G_{ij}=\mathbf{1}\Big{\{}||\alpha_{i}-\alpha_{j}||\leq\sqrt{\frac{8}{\pi N}}\Big{\}}$, where $\alpha_{i}=(\alpha_{1i},\alpha_{2i})^{\prime}$ and each component is independently drawn from a uniform distribution on $[0,1]$. This specification yields a network $\mathcal{G}_{N}$ with an expected degree of 8.
@@ B.2 Simulation Evidence
$+0.4*Y_{i}^{(t-1)}+\underbrace{0.1*\frac{\sum_{j=1}^{N}\mathbf{1}\{d_{ij}=1\}Y_{j}^{(t-1)}}{\sum_{j=1}^{N}\mathbf{1}\{d_{ij}=1\}}}_{\text{Congation}}.$
@@
For each unit $i$, its transformed outcome at each $d\in\{0,1,\dots,7\}$ is constructed as the average outcome across its $d$th-degree neighbors in the network. The propensity scores are estimated via logistic regressions fitted for each period. In each of the $1,000$ assignments, we rely on WLS to obtain $\hat{\boldsymbol{\beta}}(d)$ for each $d$ and plot them against their simulated true values in Figure B.2. Despite the moderate sample size ($400$ units), the bias is negligible, confirming the consistency of the proposed estimator.
Figure B.2: Bias of the Proposed Method
@@
We further investigate how the performance of the method changes with an increasing sample size: $N\in\{400,625,900,1225,1600\}$. The top row of Figure B.3 shows the average mean squared error (MSE) of the estimates for each $d$ and $N$, while the bottom row displays the coverage rates. As the plots indicate, MSEs decline across all proximity levels as $N$ grows, and the coverage rates approach or exceed the nominal 95% level when the sample size is sufficiently large.
Figure B.3: Asymptotic Performance of the Proposed Method
@@
B.3 Extra Results from the Applications
To assess the validity of our identification assumptions in the two applications, we conduct placebo tests as described in Section A.5, estimating the effect of treatment history $(0,1)$ on outcomes in period 00. The main analysis assumes sequential exchangeability conditional on each unit’s treatment and outcomes from the previous period, implying that outcomes in period 00 should not be influenced by the treatment status in period $2$. The results are presented in Figure B.4. Consistent with our expectation, there is no significant effect on any of the four outcomes studied in the main text, thereby supporting the identification assumptions.
Figure B.4: Placebo Tests
@@ B.3 Extra Results from the Applications
Next, we examine the robustness of our findings to potential diffusion in treatment by modifying the propensity score model to incorporate the treatment statuses and time-varying confounders of nearby neighbors. In the replication of Stokes (2016), we include constituencies within a 5 km radius. For the FHS analysis, we consider up to three degrees of neighbors. Variables from each neighbor $j$ are weighted by the inverse of $d_{ij}$ when estimating the propensity score for unit $i$. Results are presented in Figure B.5. Most point estimates and their confidence intervals remain nearly identical to those obtained without adjusting for treatment diffusion.
Figure B.5: Results with Diffusion in Treatment
@@
Finally, we report effect estimates for three additional outcomes in the FHS: blood pressure, cholesterol level, and height. Among them, height serves as a placebo outcome, as it is unlikely to be influenced by smoking cessation. The results, presented in Figure B.6, suggest that smoking cessation leads to a moderate reduction in an individual’s cholesterol level but has little effect on their social network neighbors. In contrast, neither blood pressure nor height appears to be significantly affected by the treatment.
Figure B.6: mpacts of Smoking Cessation on Additional Health Outcomes

### v5

@@ B.3 Proof for Theorem 2
Kojevnikov, Marmer and Song (2021) show that $\tilde{\Sigma}_{Y}(d)\to\Sigma_{Y}(d)$ as $N\to\infty$. Let $\mathbf{R}(d)=\frac{1}{N}\sum_{i=1}^{N}\sum_{j:\{d_{ij}\leq\bar{d}_{N}\}}{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yi}(d)\right]{\rm E}\,\left[\boldsymbol{\varepsilon}_{Yj}^{\prime}(d)\right]$. Then, by Lemma B.2,
@@
Appendix C Extra Evidence from Simulation and Applications
In this section, we present the results from our simulation study along with additional findings from the two empirical applications.
C.1 Examples of $d_{ij}$ and $\Omega_{j}(d)$
Table C.1: Examples of $d_{ij}$ and $\Omega_{j}(d)$
@@ C.1 Examples of $d_{ij}$ and $\Omega_{j}(d)$
$d_{ij}=\begin{cases}1,&\text{same street/block}\\ 2,&\text{same district/town}\\ 3,&\text{otherwise}\end{cases}$ $\{i\in\mathcal{N}:i\text{ belongs to }j\text{'s block or cluster}\}$
@@
C.2 Simulation Evidence
[…]
@@ C.2 Simulation Evidence
[…]
@@
For each unit $i$, its transformed outcome at each $d\in\{0,1,\dots,7\}$ is constructed as the average outcome across its $d$th-degree neighbors in the network. The propensity scores are estimated via logistic regressions fitted for each period. In each of the $1,000$ assignments, we rely on WLS to obtain $\hat{\boldsymbol{\beta}}(d)$ for each $d$ and plot them against their simulated true values in Figure C.2. Despite the moderate sample size ($400$ units), the bias is negligible, confirming the consistency of the proposed estimator.
Figure C.2: Bias of the Proposed Method
@@
We further investigate how the performance of the method changes with an increasing sample size: $N\in\{400,625,900,1225,1600\}$. The top row of Figure C.3 shows the average mean squared error (MSE) of the estimates for each $d$ and $N$, while the bottom row displays the coverage rates. As the plots indicate, MSEs decline across all proximity levels as $N$ grows, and the coverage rates approach or exceed the nominal 95% level when the sample size is sufficiently large.
Figure C.3: Asymptotic Performance of the Proposed Method
@@
C.3 Extra Results from the Applications
To assess the validity of our identification assumptions in the two applications, we conduct placebo tests as described in Section A.3, estimating the effect of treatment history $(0,1)$ on outcomes in period 0. The main analysis assumes sequential exchangeability conditional on each unit’s treatment and outcomes from the previous period, implying that outcomes in period 0 should not be influenced by the treatment status in period $2$. The results are presented in Figure C.4. Consistent with our expectation, there is no significant effect on any of the four outcomes studied in the main text, thereby supporting the identification assumptions.
Figure C.4: Placebo Tests
@@ C.3 Extra Results from the Applications
Next, we examine the robustness of our findings to potential diffusion in treatment by modifying the propensity score model to incorporate the treatment statuses and time-varying confounders of nearby neighbors. In the replication of Stokes (2016), we include constituencies within a 5 km radius. For the FHS analysis, we consider up to three degrees of neighbors. Variables from each neighbor $j$ are weighted by the inverse of $d_{ij}$ when estimating the propensity score for unit $i$. Results are presented in Figure C.5. Most point estimates and their confidence intervals remain nearly identical to those obtained without adjusting for treatment diffusion.
Figure C.5: Results with Diffusion in Treatment
@@
Finally, we report effect estimates for three additional outcomes in the FHS: blood pressure, cholesterol level, and height. Among them, height serves as a placebo outcome, as it is unlikely to be influenced by smoking cessation. The results, presented in Figure C.6, suggest that smoking cessation leads to a moderate reduction in an individual’s cholesterol level but has little effect on their social network neighbors. In contrast, neither blood pressure nor height appears to be significantly affected by the treatment.
Figure C.6: mpacts of Smoking Cessation on Additional Health Outcomes